OpenStudy (anonymous):

1. 2(7x – 6) = 16 x = 1.57 x = 2 x = .71 x = –2 2. 8(c + 3) –4c = –4 c = –7 c = 5 c = 7 c = –5 3. The first step to solve the following equation, –3x + 5 = –7, is: Add 5 to each side of the equation Subtract 5 from each side of the equation Divide each side of the equation by 3 Multiply each side of the equation by 3 4. Find the simple interest on $750 deposited at an interest rate of 3.5% for 3 years. $828.75 $262.50 $67.50 $78.75 5. Find the balance on $6000 deposited at 6% compounded semi-annually for 3 years. $7,164.31 $7,146.10 $8,511.11 $6,556.36 help i

OpenStudy (anonymous):

please help im helping my kid!

OpenStudy (anonymous):

@hero @Luis_Rivera @LizzyLove<3 @Katlyn99 plase maybe one of yall can help me!

OpenStudy (anonymous):

3) Subtract 5 from each side of the equation Divide each side of the equation by 3

OpenStudy (anonymous):

1) x =4, whitch i not in your options

OpenStudy (anonymous):

need help plase

OpenStudy (anonymous):

1. x=4 2. c=-7 3. x=4 4. $78.75 5. 7164.31

OpenStudy (whpalmer4):

@Luis_Rivera you read the problem wrong or something. 2(7x-6) =16 is not solved by x = 4! 2(7(4)-6) = 2(28-6) = 2*22=44 != 16 The correct answer *is* in the list of choice for problem 1.

OpenStudy (anonymous):

oh, boy !!

OpenStudy (anonymous):

sorry 1. x=2

OpenStudy (anonymous):

thank you all!